# Implication of nano-Hertz stochastic gravitational wave background on ultralight axion particles

### v4

@@ 1 Introduction
Recently, the expected Hellings-Downs [1] curve has finally been observed by the global pulsar timing array (PTA) experiments including NANOGrav [2], European PTA (EPTA) [3], Parkes PTA (PPTA) [4], and Chinese PTA (CPTA) [5], which provide strong evidence for the detection of stochastic gravitational wave background (SGWB). The observation of SGWB has wide and deep implications for astrophysics, particle physics, and cosmology. Especially, the PTA measurement of SGWB opens new windows to explore the new physics beyond the standard model [6, 7]. In this work, we focus on the implication of the new observation to the ultralight axion [8, 9, 10, 11, 12, 13, 14, 15] or axion-like particles [16] (in this letter, we simply use axion to represent both axion and axion-like particle), which could provide a natural solution to the strong CP problem and a well-motivated dark matter (DM) candidate [17, 18]. Motivated by the current situation of DM search, recently more and more studies have focused on ultralight bosons [19], such as ultralight axions. As pseudo scalar particles, ultralight axion particles can form Bose-Einstein condensation and further form dense axion objects, like axion stars or axion clusters. Especially, the so-called axion cloud can be formed through the superradiant process around Kerr black hole (BH) [20, 21, 22, 23]. Once the axion cloud is formed, abundant signals could be produced. For example, the axion cloud can be a promising gravitational wave (GW) source [24, 25, 26, 27, 28, 29, 30], which can be used to explore the properties of axion particles [31, 32, 33, 34, 35].
@@ 5 Conclusion
The assumptions we made in this work are the spin and mass distribution of SMBHs and the existence of ultralight aixon particles around $10^{-21}$ eV, which is around the mass range of the well-motivated fuzzy DM. Compared to the well-known interpretations like cosmic string, domain wall, and phase transition [6, 7] , it is less restricted by cosmological constraints and fine-tuning problems since it is formed after the Big Bang nucleosynthesis period. Our work also shows that the PTA experiments can be a novel and realistic approach to probing the superradiance or Penrose process, the isolated BHs, the ultralight axion particles or the fuzzy DM. As the leading-order term of the overlap reduction function, the Hellings-Downs curve represents the isotropic information, while the anisotropy analysis based on the higher-order terms of the overlap reduction function at future PTA experiments and its deeper implication for the sources properties (such as the detailed distribution information of BHs) are left for our future study [49].

### v5

@@ 1 Introduction
Recently, the expected Hellings-Downs [1] curve has finally been observed by the global pulsar timing array (PTA) experiments including NANOGrav [2], European PTA (EPTA) [3], Parkes PTA (PPTA) [4], and Chinese PTA (CPTA) [5], which provide strong evidence for the detection of stochastic gravitational wave background (SGWB). The observation of SGWB has wide and deep implications for astrophysics, particle physics, and cosmology. Especially, the PTA measurement of SGWB opens new windows to explore the new physics beyond the standard model [6, 7]. In this work, we focus on the implication of the new observation to the ultralight axion [8, 9, 10, 11, 12, 13, 14, 15] or axion-like particles [16] (in this letter, we simply use axion to represent both axion and axion-like particle), which could provide a natural solution to the strong CP problem and a well-motivated dark matter (DM) candidate [17, 18, 19, 20]. Motivated by the current situation of DM search, recently more and more studies have focused on ultralight bosons [21], such as ultralight axions. As pseudo scalar particles, ultralight axion particles can form Bose-Einstein condensation and further form dense axion objects, like axion stars or axion clusters. Especially, the so-called axion cloud can be formed through the superradiant process around Kerr black hole (BH) [22, 23, 24, 25]. Once the axion cloud is formed, abundant signals could be produced. For example, the axion cloud can be a promising gravitational wave (GW) source [26, 27, 28, 29, 30, 31, 32], which can be used to explore the properties of axion particles [33, 34, 35, 36, 37].
@@ 5 Conclusion
The assumptions we made in this work are the spin and mass distribution of SMBHs and the existence of ultralight aixon particles around $10^{-21}$ eV, which is around the mass range of the the well-motivated fuzzy DM. Compared to the well-known interpretations like cosmic string, domain wall, and phase transition [6, 7] , it is less restricted by cosmological constraints and fine-tuning problems since it is formed after the Big Bang nucleosynthesis period. Our work also shows that the PTA experiments can be a novel and realistic approach to probing the superradiance or Penrose process, the isolated BHs, the ultralight axion particles or the fuzzy DM. As the leading-order term of the overlap reduction function, the Hellings-Downs curve represents the isotropic information, while the anisotropy analysis based on the higher-order terms of the overlap reduction function at future PTA experiments and its deeper implication for the sources properties (such as the detailed distribution information of BHs) are left for our future study [51].